OpenStudy (anonymous):

Please help! Given the function f(x) = The quantity of 2x minus 5, divided by 3, which of the below expressions is correct?

OpenStudy (anonymous):

f−1(x) = The quantity of 5 minus 2x, divided by 3. f−1(x) = The quantity of 3x minus 5, divided by 2. f−1(x) = The quantity of 3x plus 5, divided by 2. f−1(x) = The quantity of negative 2x minus 5, divided by 3.